OpenStudy (anonymous):

Solve the following equation. u+2=4 WILL GIVE MEDAL FOR CORRECT ANSWER!!!

OpenStudy (anonymous):

Are there choices...? Because that would just be u = 2 for 2 + 2 = 4.

OpenStudy (anonymous):

No choices

OpenStudy (anonymous):

u=4-2 u=2

OpenStudy (anonymous):

Is she right?

OpenStudy (anonymous):

yeah, she's right. To equal 4 with a hidden number and 2, just add it against itself. Won't work for everything though.

OpenStudy (anonymous):

Or do what she did and take the answer and subtract it (division works too) against the other given number.

OpenStudy (anonymous):

Sorry fi that's a bit confusing.